# DHR: Dual Features-Driven Hierarchical Rebalancing in Inter- and Intra-Class Regions for Weakly-Supervised Semantic Segmentation

### v1

@@ 1 Introduction
We observe that unsupervised semantic segmentation (USS) features, as highlighted in studies [19, 33], are adept at distinguishing between inter-class regions (e.g., animal and furniture). For example, Figure 3(a) showcases pixel-wise cosine similarity maps (i.e., heatmaps) comparing WSS and USS features with class points, where USS features effectively identify differences between inter-class regions (e.g., person and motorbike). However, USS features failed to differentiate between intra-class regions (e.g., cat and dog), attributed to the absence of specific image tags during USS training. Contrarily, WSS features proficiently separate intra-class regions, as depicted in Figure 3(b), thanks to training image-level class labels.
Figure 3: Visualization of heatmaps with target points. (a): USS features can precisely separate between inter-class regions (e.g., animal and vehicle) unlike WSS. (b): Thanks to training image-level class labels, WSS features can discern specific classes (e.g., dog and cat) in the same inter-class region (e.g., animal).
@@ 2.1 Weakly-supervised Semantic Segmentation
Weakly-supervised semantic segmentation (WSS) aims to minimize the quality gap between pixel-wise annotations and pseudo masks generated with image-level class labels. Unlike previous WSS studies [38, 64, 81, 49, 14, 73, 34, 30], our research pioneers in addressing the problem of vanishing minor classes by rebalancing class information in pseudo masks in Table 1.
@@ 2.1.3 Integration with External Models
Current state-of-the-art approaches [49, 14, 73] rely on CLIP [59], large-language models [75], or SAM [35] to fine-tune WSS models. However, the dependency on pre-trained knowledge limits applicability to novel tasks, such as in medical fields, constraining WSS’s scalability. Nevertheless, our approach using both unsupervised and weakly-supervised features outperforms recent WSS methods depending on advanced supervision and datasets, demonstrating the potential of integrating WSS with USS in overcoming limitations presented by external models.
@@ 3.2.1 Step 1: Optimal Transport-based Seed Initialization
To restore minor-class regions that vanish during the traditional propagation process, we revisit CAMs $\mathcal{A}_{F^{ws}}$ utilized before the propagation $\mathcal{P}^{base}$. However, despite the presence of the small-scale class within adjacent regions in CAMs, they often overlap with larger object areas due to low-resolution CAM seeds. Consequently, in Figure 6, we first apply optimal transport (OT) to retain confident areas between CAM seeds, effectively yielding reliable class seeds akin to semantic points:
@@
Figure 6: Illustration of recovering minor-class regions from CAMs with OT. Without refining CAM seeds using OT, the result missing most minor-class areas (i.e., person) is produced. By contrast, eliminating overlapping areas within CAM seeds through OT makes it possible to restore vanished classes in the green box.
@@ 3.2.2 Step 2: USS Feature-based Rebalancing For Inter-class Regions
where $S^{us}_{ij}:=ReLU(sim(F^{us}_{ij},V^{us}))$ in $\mathbb{R}^{C}$ is the result of updating the initial mask $M^{init}$ by clustering around USS centroids $V^{us}$, aiding in the distinct categorization of inter-class regions (e.g., kitchenware and furniture). $\textit{sim}(\cdot)$ is the cosine similarity. The refined outcome, $\hat{S}^{us}$, is the result of applying OT-based optimization to this updated mask. The details of CAP are described in Appendix 0.A.2.
@@ 3.2.3 Step 3: WSS Feature-based Rebalancing For Intra-class Regions
Contrary to former approaches like $\mathcal{R}_{C}(\mathcal{P}^{base}(\mathcal{A}_{F^{ws}}))$ in (1)) which apply flawed propagation $\mathcal{P}^{base}$, our DHR initiates a new propagation, $\mathcal{P}^{ours}$, incorporating dual feature maps from USS (Step 2) and WSS (Step 3) to discern inter- and intra-class regions for all class rebalancing. We add the specific example related to the pruning operator $\mathbbm{1}_{[{sim}(V^{us},V^{us})>\tau]}$ in Appendix 0.A.3.
@@ 4.2 Comparison with State-of-the-art Approaches
DHR (Ours) ResNet-101 $\mathcal{I}$ 79.6111http://host.robots.ox.ac.uk:8080/anonymous/A4RUI9.html 79.8  [^2] http://host.robots.ox.ac.uk:8080/anonymous/HICQUU.html 53.9 49.0 32.9 37.4
@@ Appendix 0.A Method Details
0.A.1 Effect of Optimal Transport
Figure 10, an expanded illustration of Figure 6, showcases heatmaps derived from Class Activation Maps (CAMs) both with and without applying optimal transport (OT). For generating distinct and non-overlapping seeds, we use the optimal transition operator $f_{OT}(\mathcal{A}_{F^{ws}})$ to the initial CAMs $\mathcal{A}_{F^{ws}}$ before using the boundary correction tool $\mathcal{R}_{C}(\cdot)$ (i.e., CRF [37]).
Figure 10: In-depth visualization of optimal transport (OT).
0.A.2 Class-level Average Pooling
@@
0.A.3 Details of USS Feature Correlation Matrix
@@ 0.B.2 Per-class Performance Analysis
Tables 10 and 11 detail per-class segmentation outcomes for the PASCAL VOC dataset. Addressing the issue of minor-class disappearance in adjacent pixels leads to enhancements across all classes rather than improvements confined to specific categories.
@@ 0.B.3 Consistent Improvements on Other Datasets
In addition to the key component experiments conducted on the COCO dataset [48], as shown in Table 6, we extend our analysis to include detailed experimental results for the Context and ADE datasets [53, 78] (see Tables 7 and 8). These further analyses show that our DHR method enhances performance significantly, achieving up to a 17.9% improvement on two datasets, which feature adjacent class scenarios.
@@ 0.C.1 Model-agnostic Improvements
Figures 13 and 14 provide a qualitative comparison between our DHR method, two baseline models [68, 41], and various model-agnostic approaches [50, 42, 30]. Demonstrating its robustness, DHR excels in segmenting diverse objects and managing scenarios with multiple classes. It particularly shines in accurately segmenting adjacent minor-class regions (e.g., person and bicycle), where it significantly outperforms other WSS methods [50, 42, 30] by recovering classes that are often missed or overlooked, thereby ensuring comprehensive and satisfactory segmentation outcomes on five benchmarks.

### v2

@@ 1 Introduction
We observe that unsupervised semantic segmentation (USS) features, as highlighted in studies [22, 36], are adept at distinguishing between inter-class regions (e.g., animal and furniture). Figure 3 showcases pixel-wise cosine similarity maps (i.e., heatmaps) comparing WSS and USS features with class points. Specifically, the inherent difference between USS and WSS features originates from their distinct learning objectives. In Figure 3(a), USS methods learn visual similarities (e.g., color and shape) between images without tags, enabling them to distinguish inter-class regions with dissimilar appearances (e.g., person vs. motorbike). Conversely, WSS methods learn class-specific differences with human-annotated tags, allowing them to discern visually similar classes (e.g., dog vs. cat) within the same inter-class region (e.g., animals), as shown in Figure 3(b).
Figure 3: Visualization of heatmaps with target points. (a): USS features can precisely separate between inter-class regions (e.g., animals vs. vehicles) unlike WSS. (b): Thanks to training image-level class labels, WSS features can discern specific classes (e.g., dog vs. cat) in the same inter-class region (e.g., animal).
@@ 2.1 Weakly-supervised Semantic Segmentation
Weakly-supervised semantic segmentation (WSS) aims to minimize the quality gap between pixel-wise annotations and pseudo masks generated with image-level class labels. Unlike previous WSS studies [21, 11, 72, 9, 41, 68, 86, 52, 16, 78, 37, 33], our research pioneers in addressing the problem of vanishing minor classes by rebalancing class information in pseudo masks in Table 1.
@@ 2.1.3 Integration with External Models
Current state-of-the-art approaches [52, 16, 78, 9, 11, 72] rely on CLIP [63], large-language models [80], Grounding DINO [54], or SAM [38] to fine-tune WSS models. However, the dependency on pre-trained knowledge limits applicability to novel tasks, such as in medical fields, constraining WSS’s scalability. Nevertheless, our approach using both unsupervised and weakly-supervised features outperforms recent WSS methods depending on advanced supervision and datasets, demonstrating the potential of integrating WSS with USS in overcoming limitations presented by external models.
@@ 3.2.1 Step 1: Optimal Transport-based Seed Initialization
To restore minor-class regions that vanish during the traditional propagation process, we revisit CAMs $\mathcal{A}_{F^{ws}}$ utilized before the propagation $\mathcal{P}^{base}$. As shown in Figure 6(a), we first observe that Optimal Transport (OT) [62] effectively minimizes the occurrence of false positives (FP) in regions adjacent to the target classes:
@@
Figure 6: Illustration of recovering minor-class regions from CAMs with OT. (a): OT reduces false positives in adjacent class regions (e.g., bottle vs. bicycle). (b): The input WSS mask often misses most minor-class areas (i.e., person) in the red box. By contrast, refining CAM seeds with OT eliminates overlapping areas, restoring minor-class regions that have previously vanished, as indicated in the green box.
@@ 3.2.2 Step 2: USS Feature-based Rebalancing For Inter-class Regions
where $S^{us}_{ij}:=ReLU(sim(F^{us}_{ij},V^{us}))$ in $\mathbb{R}^{C}$ is the result of updating the initial mask $M^{init}$ by clustering around USS centroids $V^{us}$, aiding in the distinct categorization of inter-class regions (e.g., kitchenware and furniture). $\textit{sim}(\cdot)$ is the cosine similarity. The refined outcome, $\hat{S}^{us}$, is the result of applying OT-based optimization to this updated mask. We describe CAP in Appendix 0.A.1.
@@ 3.2.3 Step 3: WSS Feature-based Rebalancing For Intra-class Regions
Contrary to former approaches like $\mathcal{R}_{C}(\mathcal{P}^{base}(\mathcal{A}_{F^{ws}}))$ in (1)) which apply flawed propagation $\mathcal{P}^{base}$, our DHR initiates a new propagation, $\mathcal{P}^{ours}$, incorporating dual feature maps from USS (Step 2) and WSS (Step 3) to discern inter- and intra-class regions for all class rebalancing. We add the specific example related to the pruning operator $\mathbbm{1}_{[{sim}(V^{us},V^{us})>\tau]}$ in Appendix 0.A.2.
@@ 4.2 Comparison with State-of-the-art Approaches
SEPL arXiv’23 [11] ResNet-101 $\mathcal{I}$+$\mathcal{S}$+$\mathcal{C}$ 71.1 - - - -
@@
ROSE Information Fusion’24 [9] ResNet-101 $\mathcal{I}$+$\mathcal{S}$+$\mathcal{C}$ 75.4 76.6 48.3 - - -
Sun et al. arXiv’23 [72] ResNet-101 $\mathcal{I}$+$\mathcal{S}$+$\mathcal{D}$ 77.2 77.1 55.6 - - -
FMA-WSSS WACV’24 [78] ResNet-101 $\mathcal{I}$+$\mathcal{S}$+$\mathcal{C}$ 77.3 76.7 48.6 - - -
@@
DHR (Ours, DeepLabv3+) ResNet-101 $\mathcal{I}$ 79.6111http://host.robots.ox.ac.uk:8080/anonymous/A4RUI9.html 79.8  [^2] http://host.robots.ox.ac.uk:8080/anonymous/HICQUU.html 53.9 49.0 32.9 37.4
DHR (Ours, Mask2Former) Swin-L $\mathcal{I}$ 82.3 82.3 56.8 53.6 36.9 41.1
Upper Bound (DeepLabv3+) CVPR’18 [8] ResNet-101 $\mathcal{M}$ 80.6* 81.0* 61.8* 54.6* 45.3* 44.2*
Upper Bound (Mask2Former) CVPR’22 [13] Swin-L $\mathcal{M}$ 86.0 86.1 66.7 64.3* 55.5* 50.6*
@@ Appendix 0.A Method Details
0.A.1 Class-level Average Pooling
@@
0.A.2 Details of USS Feature Correlation Matrix
@@
0.A.3 Computational Complexity
To quantify the computational overhead of applying our DHR, we investigate the training and testing times across five datasets, as detailed in Table 7. The total training time with DHR increases by a factor of 1.8 (from 10 hours to 18 hours on VOC 2012 [18]). To mitigate the computational overhead introduced by techniques such as OT [62], we utilize 64 CPU cores in parallel, effectively reducing the impact on the overall training time. Importantly, despite the increased complexity during training, the testing time remains consistent with all baselines [1, 73, 44, 32]. This consistency ensures that our refining steps are applied exclusively during the training phase, maintaining model deployment efficiency.
Table 7: Complexity comparison of our baseline (i.e., RSEPM [32]) with and without DHR. As all DHR steps in Sec. 3.2 are used only during training, the inference time is the same as all baselines [1, 73, 44, 32].
(Dataset) Phase RSEPM [32] without DHR RSEPM [32] with DHR
(VOC 2012) Total Training Time 10 hours 18 hours (+8 hours)
(VOC 2012) Total Testing Time 5 minutes 5 minutes (+0 minutes)
(COCO 2014) Total Training Time 19 hours 36 hours (+17 hours)
(COCO 2014) Total Testing Time 30 minutes 30 minutes (+0 minutes)
(Context) Total Training Time 5 hours 9 hours (+4 hours)
(Context) Total Testing Time 10 minutes 10 minutes (+0 minutes)
(ADE 2016) Total Training Time 11 hours 20 hours (+9 hours)
(ADE 2016) Total Testing Time 15 minutes 15 minutes (+0 minutes)
(Stuff) Total Training Time 40 hours 73 hours (+33 hours)
(Stuff) Total Testing Time 20 minutes 20 minutes (+0 minutes)
@@ 0.B.2 Per-class Performance Analysis
Tables 11 and 12 detail per-class segmentation outcomes for the PASCAL VOC dataset. Addressing the issue of minor-class disappearance in adjacent pixels leads to enhancements across all classes rather than improvements confined to specific categories.
@@ 0.B.3 Consistent Improvements on Other Datasets
In addition to the key component experiments conducted on the COCO dataset [51], as shown in Table 6, we extend our analysis to include detailed experimental results for the Context and ADE datasets [56, 83] (see Tables 8 and 9). These further analyses show that our DHR method enhances performance significantly, achieving up to a 17.9% improvement on two datasets, which feature adjacent class scenarios.
@@ 0.C.1 Model-agnostic Improvements
Figures 12 and 13 provide a qualitative comparison between our DHR method, two baseline models [73, 44], and various model-agnostic approaches [53, 45, 33]. Demonstrating its robustness, DHR excels in segmenting diverse objects and managing scenarios with multiple classes. It particularly shines in accurately segmenting adjacent minor-class regions (e.g., person and bicycle), where it significantly outperforms other WSS methods [53, 45, 33] by recovering classes that are often missed or overlooked, thereby ensuring comprehensive and satisfactory segmentation outcomes on five benchmarks.